NASA ISS On-Orbit Status 21 January 2014

©NASA

NASA ISS On-Orbit Status 21 January 2014.

Today: Cosmo Shoot Operations: Flight Engineer (FE)-6 Wakata has set up the Cosmo Shoot laptop and 4K camera in the Cupola to capture images of the Earth surface. He later removed the hardware and temporarily stowed it in the Japanese Experiment Module (JEM).

ENERGY (Astronaut's Energy Requirements for Long-Term Space Flight): FE-3 Hopkins has started his standardized diet protocol for a 10 day session of ESA's ENERGY experiment. Measurements have been taken of his oxygen uptake during rest periods, urine collections will be performed and Hopkins will log all food and drink consumed. A loss in an astronauts' body mass is common during space flight. ENERGY measures changes in the energy balance in crewmembers as a result of long term space flight. It also measures adaptations in the components of total energy expenditure of crewmembers, so that an equation to determine the energy requirements of crewmembers during space flight may be derived. Such knowledge is of great importance to ensure optimal crew health and performance, overall mission success, and to ensure adequate exercise load and cargo allotments for food are provided while the crew is on-orbit.

Retractable Equipment Tether (RET) Inspection and Tool Transfer: Yesterday, FE-5 Mastracchio performed a RET inspection in preparation for today's transferring of the hardware to the Russian crewmembers for use on Russian Segment (RS) Extravehicular Activity (EVA) 37A. Overnight, engineering reviewed the pictures from the Bungee Caddies and as a result could not clear them for use with pictures alone. Ground teams are currently working on a plan to perform additional inspections on this hardware. In the meantime, in order to get the tools transferred to the RS on time, the crew inspected unmounted RETs to use in place of the mounted RETs. All unmounted RETs passed inspection and were included in the EVA tool transfer to the RS today. RS EVA 37A is currently planned for January 27th.

Sound Level Meter (SLM) Operations: Mastracchio performed SLM measurements throughout the ISS, collecting acoustic readings in JEM Pressurized Module (JPM), Cygnus, Columbus, Node 2, Lab, Cupola, Node 1, Pressurized Mating Adapter (PMA)-1, Mini Research Module (MRM)-1, and the Service Module (SM). The data collected will be downlinked and reviewed by specialists to determine if any mitigation steps are required for the various modules.

Carbon Dioxide Removal System (CDRA) Leak Check: Ground controllers completed a series of leak checks on the Node 3 CDRA as part of an ongoing investigation into the moisture observed in Sabatier CO2 accumulator. The procedure checked for leaks of the volume between each of the Adsorbent/Desiccant beds and the CO2 shutoff valve. Initial review of the data showed no gross leaks.

Today's Planned Tasks

ENERGY - PFM/PAM Activation [Complete]
Regeneration of БМП Ф2 Micropurification Cartridge (start) [Complete]
ENERGY - Urine Sample Collection [Complete]
ENERGY Instrument Check [Complete]
ENERGY Double Labeled Water Intake [Complete]
Energy Oxygen Uptake - Part 1 [Complete]
Activation and inspection of Orlan suits No. 4, No. 6 [Complete]
COSMO - HDTV Camera Setup [Complete]
Checkout of БСС (Spacesuit Interface Unit) in ПхО [Complete]
Checkout of БСС (Spacesuit Interface Unit) in DC1 [Complete]
COSMO - Camera Preparation [Complete]
ENERGY - BREAKFAST [Complete]
Orlan Suit No. 4, No. 6, and БСС Cooling Loop Separation in DC1 [Complete]
Energy Oxygen Uptake - Part 2 [Complete]
COSMO - Camera View Check [Complete]
COSMO - Camera lens exchange [Complete]
Cosmo Shoot [Complete]
Energy Urine Sample Collection [Complete]
VIRTUAL. Setup, Conduct Experiment and Closeout Ops [Complete]
Cleaning Fluid System by using Degassing Pump Unit (БОС) in ПхО [Complete]
Energy Oxygen Uptake - Part 3 [Complete]
COSMO - Camera lens exchange [Complete]
Cleaning ВД1 and ВД2 Air Ducts in MRM2 [Complete]
WRS Water Sample Analysis [Complete]
PAO Hardware Setup [Complete]
Perform RET splice measurement on additional RETs being used for Russian EVA 37A. [Complete]
Replacement of ПФ1, ПФ2 Dust Filters and Cleaning В1, B2 Fan Screens in MRM2 [Complete]
Handover of US EVA Equipment and Tools to RS Crew [Complete]
ENERGY - Urine Sample Collection [Complete]
Combustion Integrated Rack Alignment Guide Installation [Complete]
TERMO-LADA. Visual Inspection of Hardware Operation [Complete]
Energy Oxygen Uptake - Part 4 [In Work]
Microbial Monitoring of the Environment (sample collection for the first kit) [Complete]
WRM - Condensate Pumping Initiation [Complete]
ENERGY - Urine Sample Collection [Complete]
Filling (separation) of EDV [KOV] for Elektron [Complete]
WRM - Terminate Condensate Pumping [Complete]
СОЖ Maintenance [Complete]
Gather BASS samples, InSpace tapes, in a 1.0 CTB, then stage for SpX-3 return [Complete]
ENERGY Diet Log of Lunch [Complete]
ENERGY - Closeout Ops [Complete]
HOSC - Hard Drive Installation [Complete]
ENERGY - File transfer to laptop [Complete]
Sound Level Meter Ops [Complete]
ENERGY - Data Transfer and Equipment Power Down [Complete]
Bracket Stowage [Complete]
Working with space suits replaceable elements, inspection [Complete]
Transfer of 3.0 CTB CWC-I from PMM to Node 3 [Complete]
Replacement of СРВК-2М Gas-Liquid Mixture Filter (ФГС) Install ФГС No. 1009019 Discard the removed unit. [Complete]
COSMO -Camera View Check [Complete]
Replacement of СД1-7 Light in DC1 [Complete]
MOTOCARD. Experiment Ops [Complete]
MOTOCARD. Experiment Assistance [Complete]
SLM data transfer [Complete]
Terminate and remove from charging charging the 2nd Orlan-MK battery pack 825M3 [Complete]
TOCA Data Recording [Complete]
TOCA Waste Water Bag (WWB) Changeout [Complete]
ТКГ 420 Packing and IMS Ops [Complete]
RGN - Initiate drain into EDV [Complete]
IMS Update [Complete]
COSMO - Camera lens exchange [Complete]
RGN - Terminate drain into EDV [Complete]
COSMO - Camera View Check [Complete]
БМП Ф2 Absorption Cartridge Regeneration (termination) [Complete]
COSMO - Hardware Deactivation, Closeout [Complete]
COSMO - Temporary Stowage [Complete]
Preparation for МО-21. Ecosfera Battery Charge [Complete]
Cupola Window Shutter Close [Complete]
ENERGY Diet Log of Dinner [Complete]
Task List

ARED Cylinder Evac
National Geographic B Roll Part 4 [In Work]
ЕДВ-У R&R
Replace KTO when full.
Preparation of reports for Roskosmos site
URAGAN. Observation and photography
EKON-M. Observations and photography
CFE VG2 Ops
St Moon Read [Deferred]
JAXA Report
Ground

Nominal Commanding [Complete]
Cupola Window Heaters Powerdown Test [In Work]
Three-Day Look Ahead:

Wednesday, 01/22: SPHERES Slosh, N2 fwd endcone cleanout
Thursday, 01/23: OCT, Fundoscope, DRAGON prepack
Friday, 01/24: SPHERES Vertigo, Ocular Health Cardiac
QUICK ISS Status - Environmental Control Group:

Component - Status
Elektron - Off
Vozdukh - Manual
[СКВ] 1 - SM Air Conditioner System ("SKV1") - Off
[СКВ] 2 - SM Air Conditioner System ("SKV2") - Off
Carbon Dioxide Removal Assembly (CDRA) Lab - Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 - Operate
Major Constituent Analyzer (MCA) Lab - Off
Major Constituent Analyzer (MCA) Node 3 - Operate
Oxygen Generation Assembly (OGA) - Process
Urine Processing Assembly (UPA) - Standby
Trace Contaminant Control System (TCCS) Lab - Off
Trace Contaminant Control System (TCCS) Node 3 - Full up

Please follow SpaceRef on Twitter and Like us on Facebook.